OpenStudy (anonymous):

Find the x-intercept of x^4+5x^3+4x^2???? Please help :)

OpenStudy (ranga):

Factor out x^2 first. Then equate it to zero and solve for x.

OpenStudy (anonymous):

So do you mean (x^2+4x)(x^2+x)?

OpenStudy (ranga):

y = x^4+5x^3+4x^2 = x^2(x^2 + 5x + 4) Factor the quadratic: y = x^2(x+4)(x+1) x-intercept is when y = 0 x^2(x+4)(x+1) = 0 Solve for x.

OpenStudy (anonymous):

Oh. Yes it is correct thank you so much :) I though I had to do easy tri, not just simply factor it lol

OpenStudy (ranga):

you are welcome.